# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v1

@@ IV PBH abundance with non-Gaussian curvature perturbation
The non-Gaussian curvature perturbation discussed in the previous section may have a significantly different tail from the Gaussian tail in the PDF in the large $\mathcal{R}$ regime. As we commented, the quantum fluctuations of the inflaton field on subhorizon scales naturally provide a random variable $\delta\varphi$ which has a Gaussian distribution among all the Hubble patches, as the quantum fluctuation is usually Gaussian, especially when the potential is approximately quadratic. Usually, one can define the Gaussian component of the curvature perturbation $\mathcal{R}_{\mathrm{g}}$. In most of the cases, it can be defined as proportional to the inflaton field perturbation $\delta\varphi$ which recovers the leading order Gaussian relation $\mathcal{R}_{\mathrm{g}}\equiv-(H/\dot{\varphi})\delta\varphi$. As we have discussed in Section II.2, the nonlinear relation $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$ can bring non-Gaussianities to the PDF of $\mathcal{R}$, as is shown in (17). Then the PBH abundance should be calculated following (20), by transferring the PDF $\mathbb{P}(\mathcal{R})$ to that of the linear compaction function $\mathbb{P}(\mathscr{C}_{\ell})$.
For instance, in ultra-slow-roll inflation, the nonlinear relation $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$ is given by (30):
@@
$\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})=-\frac{1}{3}\ln\left(1-3\mathcal{R}_{\mathrm{g}}\right),\quad$ (60)
where $\mathrm{with}~{}\mathcal{R}_{\mathrm{g}}\equiv\delta\pi_{t}/\pi_{t}$. Then the PDF of $\mathcal{R}$, given by (18), is
@@
$\mathbb{P}(\mathcal{R})=\frac{e^{-3\mathcal{R}}}{\sqrt{2\pi}\sigma_{\mathcal{R}_{\mathrm{g}}}}\exp\left[-\frac{\left(e^{-3\mathcal{R}}-1\right)^{2}}{18\sigma^{2}_{\mathcal{R}_{\mathrm{g}}}}\right].$ (61)
@@
Unlike the Gaussian PDF, it goes as $\exp(-3\mathcal{R})$ when $\mathcal{R}\gg\mathcal{O}(0.1)$, so is called exponential tail PDF. Apprarently, for the bumpy potential (constant-roll) case, we have a similar expression with $\mathbb{P}(\mathcal{R})\propto\exp(\lambda_{-}\mathcal{R})$. Such exponential tails are less suppressed compared to the Gaussian tails, thus the PBH abundance will be greatly enhanced. Although $\mathbb{P}(\mathcal{R})$ is not directly used in the calculation of PBH mass function, as we use $\mathbb{P}(\mathscr{C}_{\ell})$ in the Press-Schechter-type formalism, it is straightforward to see the origin of such enhancement by the tail of $\mathbb{P}(\mathcal{R})$ when $\mathcal{R}\gg\sigma_{\mathcal{R}_{\mathrm{g}}}$.
@@
We will show how to calculate the PBH abundance for non-Gaussian curvature perturbation (60) in the extended Press-Schechter formalism, following Gow et al. (2022). By using (19), The compaction function $\mathscr{C}_{\ell}$ can be written as
@@
$\mathscr{C}_{\ell}\equiv\frac{X}{Y},\qquad\mathrm{with}\quad X\equiv-\frac{4}{3}r\mathcal{R}^{\prime}_{\mathrm{g}}(r),\quad Y\equiv 1-3\mathcal{R}_{\mathrm{g}}.$ (62)
@@
The covariance matrix is $\boldsymbol{\Sigma}=\left(\begin{matrix}\sigma^{2}_{X}&\sigma^{2}_{XY}\\ \sigma^{2}_{YX}&\sigma^{2}_{Y}\end{matrix}\right)$, with
@@
$\sigma^{2}_{X}$ $=\left(-\frac{4}{3}\right)^{2}\int(kr)^{2}\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)\left(\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(kr)\right)^{2}\mathrm{d}\ln k,$ (63)
@@
$\sigma^{2}_{XY}=\sigma^{2}_{YX}$ $=\left(-\frac{4}{3}\right)(-3)\int(kr)\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)j_{0}(kr)\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(kr)\mathrm{d}\ln k,$ (64)
$\sigma^{2}_{Y}$ $=(-3)^{2}\int\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)j_{0}^{2}(kr)\mathrm{d}\ln k,$ (65)
where $j_{0}(x)=\sin z/z$ is the 0th order spherical Bessel function. The joint PDF of $X$ and $Y$ becomes
@@
$\mathbb{P}(X,Y)$ $=\frac{1}{2\pi\sqrt{\mathrm{det}\boldsymbol{\Sigma}}}\exp\left[-\frac{1}{2}(X,1-Y)\boldsymbol{\Sigma}^{-1}\left(\begin{matrix}X\\ 1-Y\end{matrix}\right)\right]$ (68)
$=\frac{1}{2\pi\sigma_{X}\sigma_{Y}\sqrt{1-\varrho^{2}}}\exp\left(-\frac{X^{2}}{2\sigma^{2}_{X}}\right)\exp\left[-\frac{1}{2\left(1-\varrho^{2}\right)}\left(\frac{1-Y}{\sigma_{Y}}-\frac{\varrho}{\sigma_{X}}X\right)^{2}\right],$ (69)
where $\varrho\equiv\sigma_{XY}^{2}/(\sigma_{X}\sigma_{Y})$ is the correlation. Then the PDF of $\mathscr{C}_{\ell}$ reads
@@
$\mathbb{P}(\mathscr{C}_{\ell})=\int\mathrm{d}X\mathrm{d}Y~{}\mathbb{P}(X,Y)\delta\left(\mathscr{C}_{\ell}-\frac{X}{Y}\right)=\int\mathrm{d}X~{}Y\cdot\mathbb{P}\left(X\to\mathscr{C}_{\ell}Y,Y\right).$ (70)
@@
Another solution with $\mathscr{C}_{\ell}>4/3$ is discarded as it is the Type II PBH. With (71), to change the variable from $\mathscr{C}_{\ell}$ to $M$, we have
@@
The PBH abundance today $f_{\mathrm{PBH}}=\int f_{\mathrm{PBH}}(M)\mathrm{d}\ln M$ over all the mass range, which can be used to determine $\mathcal{P_{R}}$ needed. For PBH as all the dark matter, $f_{\mathrm{PBH}}=1$.
As an example, let’s study the monochromatic power spectrum
@@
$\sigma_{X}=\frac{4}{3}(k_{*}r)\mathcal{A}_{\mathcal{R}}^{1/2}\left|\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(k_{*}r)\right|,\qquad\sigma_{Y}=3\mathcal{A}_{\mathcal{R}}^{1/2}\left|j_{0}(k_{*}r)\right|,$ (77)
@@
and $\varrho=1$. For simplicity, we choose $r_{m}$ to be the same as the Gaussian case, $k_{*}r_{m}\approx 2.74$. Therefore, the PDF (70) degenerate into one-dimensional, as $X$ and $Y$ are fully correlated by $Y=1-(\sigma_{Y}/\sigma_{X})X$, which, together with $X=\mathscr{C}_{\ell}Y$, gives
$X=\frac{\sigma_{X}\mathscr{C}_{\ell}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}},\qquad Y=\frac{\sigma_{X}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}},\qquad$ (78)
And the degenerate PDF of $\mathscr{C}_{\ell}$ becomes
@@
$\mathbb{P}(\mathscr{C}_{\ell})=\frac{1}{\sqrt{2\pi\sigma^{2}_{X}}}\left(\frac{\sigma_{X}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}}\right)^{2}\exp\left[-\frac{1}{2\sigma^{2}_{X}}\left(\frac{\sigma_{X}\mathscr{C}_{\ell}}{\sigma_{X}+\sigma_{Y}\mathscr{C}_{\ell}}\right)^{2}\right].$ (79)
@@
We calculated the PBH mass function $f_{\mathrm{PBH}}$ for a monochromatic power spectrum in the ultra-slow-roll inflation, shown in the left panel of Fig. 6. The amplitude of the power spectrum of the curvature perturbation $\mathcal{A_{R}}$ is set to $4.151077\times 10^{-3}$, such that $f_{\mathrm{PBH}}=1$, while the $f_{\mathrm{PBH}}(M)$ peaks around $10^{-12}$ solar mass, right in the “asteroid mass window” where PBHs can be all the dark matter. Interestingly, as is clearly seen in Fig. 6, for the narrow peak case, the universal infrared scaling $f_{\mathrm{PBH}}(M)\propto M^{1+1/\gamma}\sim M^{3.78}$ which originates from the critical behavior is sharper than the constraints from the extra-galactic $\gamma$ rays proportional to $M^{3.1\sim 3.4}$ Carr et al. (2010). So the entire asteroid-mass window is open for narrow-peak power spectrum even with non-Gaussianity.
@@
By the same method, for perturbative series of (21), we can calculate the PBH abundance $f_{\mathrm{PBH}}$ as a function of $\mathcal{A_{R}}$ for different $f_{\mathrm{NL}}$’s. This is shown in the right panel of Fig. 6 together with the ultra-slow-roll inflation. We see clearly that a positive $f_{\mathrm{NL}}$ systematically enhance the PBH abundance, which makes the $\mathcal{A_{R}}$ required to realize $f_{\mathrm{PBH}}=1$ smaller, as we commented in the introduction. Also, because of the higher order terms, ultra-slow-roll inflation with fully nonlinear relation (60) requires slightly larger power spectrum compared to the quadratic expansion, although both of them have $f_{\mathrm{NL}}=5/2$. The $\mathcal{A_{R}}$ required to realize PBH dark matter for $M_{k_{*}}=2\times 10^{21}~{}\text{g}$ are listed in Table 1.
@@
Figure 6: Left: The PBH mass function $f_{\mathrm{PBH}}(M)$ given by (75) with a monochromatic power spectrum (76) in a ultra-slow-roll inflation with a sharp end, together with the observational constraints from Carr et al. (2020). The total power and central mass are $\mathcal{A_{R}}=4.151\times 10^{-3}$ and $M_{k_{*}}=2\times 10^{21}~{}\text{g}$, such that $\int f_{\mathrm{PBH}}(M)\mathrm{d}\ln M=1$. Right: The PBH abundance $f_{\mathrm{PBH}}$ as a function of the total power $\mathcal{A_{R}}$, for a monochromatic power spectrum (76). From right to left, the black, purple dotted, blue dashed, cyan, brown dot-dashed, and gray curves are for Gaussian case, $f_{\mathrm{NL}}=1$, $f_{\mathrm{NL}}=5/2$, ultra-slow-roll inflation, $f_{\mathrm{NL}}=10$, $f_{\mathrm{NL}}=10^{2}$, and $f_{\mathrm{NL}}=10^{3}$, respectively.
@@
1 $4.804923\times 10^{-3}$
5/2 $4.010815\times 10^{-3}$
5/2 (USR) $4.151077\times 10^{-3}$
10 $2.299748\times 10^{-3}$
50 $8.051903\times 10^{-4}$
$10^{2}$ $4.674072\times 10^{-4}$
$10^{3}$ $6.268813\times 10^{-5}$
Table 1: The total power $\mathcal{A_{R}}$ for a monochromatic power spectrum (76) needed to make PBH to be all the dark matter at $M_{k_{*}}=2\times 10^{21}~{}\text{g}$, with different $f_{\mathrm{NL}}$’s. The fourth line is the ultra-slow-roll model with a logarithmic $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ relation (60), although it also has $f_{\mathrm{NL}}=5/2$.

### v2

@@ IV PBH abundance with non-Gaussian curvature perturbation
The non-Gaussian curvature perturbation discussed in the previous section may have a significantly different tail from the Gaussian tail in the PDF in the large $\mathcal{R}$ regime. As we commented, the quantum fluctuations of the inflaton field on subhorizon scales naturally provide a random variable $\delta\varphi$ which has a Gaussian distribution among all the Hubble patches, as the quantum fluctuation is Gaussian when the potential is approximately quadratic. In most of the cases, the Gaussian component of the curvature perturbation $\mathcal{R}_{\mathrm{g}}$ can be defined as proportional to the inflaton field perturbation $\delta\varphi$ which recovers the leading order Gaussian relation $\mathcal{R}_{\mathrm{g}}\equiv-(H/\dot{\varphi})\delta\varphi$. As we have discussed in Section II.2, the nonlinear relation $\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})$ can bring non-Gaussianities to the PDF of $\mathcal{R}$, as is shown in (17). Then the PBH abundance should be calculated following (20), by transferring the PDF $\mathbb{P}(\mathcal{R})$ to that of the linear compaction function $\mathbb{P}(\mathscr{C}_{\ell})$.
For instance, consider the following logarithmic relation of $\mathcal{R}$,
@@
$\mathcal{R}=\mathscr{F}(\mathcal{R}_{\mathrm{g}})=\frac{1}{\lambda}\ln\left(1+\lambda\mathcal{R}_{\mathrm{g}}\right),\quad$ (57)
where $\mathcal{R}_{\mathrm{g}}$ is Gaussian. This formula represents ultra-slow-roll inflation ($\lambda=-3$), constant-roll inflation ($\lambda=\lambda_{-}$ defined in (34)), and $r\to 1$ limit of the curvaton scenario ($\lambda=3/2$). Knowing the PDF of $\mathcal{R}_{\mathrm{g}}$ and the Jacobian (18), the PDF of $\mathcal{R}$ is given by
@@
$\mathbb{P}(\mathcal{R})=\frac{e^{\lambda\mathcal{R}}}{\sqrt{2\pi}\sigma_{\mathcal{R}\mathrm{g}}}\exp\left[-\frac{\left(e^{\lambda\mathcal{R}}-1\right)^{2}}{2\lambda^{2}\sigma_{\mathcal{R}\mathrm{g}}^{2}}\right],$ (58)
@@
where $\sigma_{\mathcal{R}\mathrm{g}}$ is the root mean square of the Gaussian perturbation $\mathcal{R}_{\mathrm{g}}$, defined in (16). For a negative $\lambda$ (including $\lambda=-3$), it goes as $\exp(\lambda\mathcal{R})$ when $\mathcal{R}\gg\mathcal{O}(0.1)$, so is called exponential tail PDF. Such a PDF is less suppressed compared to the Gaussian tail, thus the PBH abundance will be greatly enhanced. Although $\mathbb{P}(\mathcal{R})$ is not directly used in the calculation of PBH mass function in the Press-Schechter-type formalism, it is straightforward to see the origin of such enhancement by the tail of $\mathbb{P}(\mathcal{R})$ when $\mathcal{R}\gg\sigma_{\mathcal{R}_{\mathrm{g}}}$. On the contrary, for a positive $\lambda$, the double exponential suppression is dominant, $\mathbb{P}\propto\exp(-c^{2}e^{\lambda\mathcal{R}})$, called Gumbel-distribution-like PDF.
@@
We will show how to calculate the PBH abundance for non-Gaussian curvature perturbation (57) in the Press-Schechter-type formalism, following Gow et al. (2022). By using (19), the linear compaction function $\mathscr{C}_{\ell}$ can be written as
@@
$\mathscr{C}_{\ell}\equiv\frac{X}{Y},\qquad\mathrm{with}\quad X\equiv-\frac{4}{3}r\mathcal{R}^{\prime}_{\mathrm{g}}(r),\quad Y\equiv 1+\lambda\mathcal{R}_{\mathrm{g}}.$ (59)
@@
The covariance matrix is $\boldsymbol{\Sigma}=\left(\begin{matrix}\sigma^{2}_{X}&\varrho\sigma_{X}\sigma_{Y}\\ \varrho\sigma_{X}\sigma_{Y}&\sigma^{2}_{Y}\end{matrix}\right)$, with
@@
$\sigma^{2}_{X}$ $=\langle XX\rangle=\left(-\frac{4}{3}\right)^{2}\int(kr)^{2}\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)\left(\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(kr)\right)^{2}\mathrm{d}\ln k,$ (60)
@@
$\varrho\sigma_{X}\sigma_{Y}$ $=\langle XY\rangle=\left(-\frac{4}{3}\right)\lambda\int(kr)\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)j_{0}(kr)\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(kr)\mathrm{d}\ln k,$ (61)
$\sigma^{2}_{Y}$ $=\langle YY\rangle=\lambda^{2}\int\mathcal{P}_{\mathcal{R}\mathrm{g}}(k)j_{0}^{2}(kr)\mathrm{d}\ln k,$ (62)
where $\rho$ is the correlation between $X$ and $Y$, $j_{0}(z)=\sin z/z$ is the zeroth order spherical Bessel function, and we assume a spherical $\zeta$ profile of
$\mathcal{R}_{\mathrm{g}}(r)=\frac{1}{\sqrt{2}\pi}\int\mathrm{d}k~{}kj_{0}(kr)\mathcal{R}_{k},\qquad\langle\mathcal{R}_{k}\mathcal{R}_{p}\rangle=\delta(k-p)\frac{2\pi^{2}}{k^{3}}\mathcal{P}_{\mathcal{R}\mathrm{g}}(k).$ (63)
The joint PDF of $X$ and $Y$ becomes
@@
$\mathbb{P}(X,Y)$ $=\frac{1}{2\pi\sqrt{\mathrm{det}\boldsymbol{\Sigma}}}\exp\left[-\frac{1}{2}(X,\,Y-1)\boldsymbol{\Sigma}^{-1}\left(\begin{matrix}X\\ Y-1\end{matrix}\right)\right]$
$=\frac{1}{2\pi\sigma_{X}\sigma_{Y}\sqrt{1-\varrho^{2}}}\exp\left(-\frac{X^{2}}{2\sigma^{2}_{X}}\right)\exp\left[-\frac{1}{2\left(1-\varrho^{2}\right)}\left(\varrho\frac{X}{\sigma_{X}}-\frac{Y-1}{\sigma_{Y}}\right)^{2}\right],$ (64)
which gives the PDF of $\mathscr{C}_{\ell}$,
@@
$\mathbb{P}(\mathscr{C}_{\ell})=\int\mathrm{d}X\mathrm{d}Y~{}\mathbb{P}(X,Y)\delta\left(\mathscr{C}_{\ell}-\frac{X}{Y}\right)=\int\mathrm{d}X~{}|Y|\cdot\mathbb{P}\left(X\to\mathscr{C}_{\ell}Y,Y\right).$ (65)
@@
Another solution with $\mathscr{C}_{\ell}>4/3$ is discarded as it is for the Type II PBH. With (66), to change the variable from $\mathscr{C}_{\ell}$ to $M$, we have
@@
The PBH abundance today $f_{\mathrm{PBH}}=\int f_{\mathrm{PBH}}(M)\mathrm{d}\ln M$, is the integration of the mass function over all the mass range. For PBH as all the dark matter, $f_{\mathrm{PBH}}=1$.
For simplicity, we only consider the monochromatic power spectrum
@@
$\sigma_{X}=\frac{4}{3}(k_{*}r)\mathcal{A}_{\mathcal{R}}^{1/2}\left|\frac{\mathrm{d}j_{0}}{\mathrm{d}z}(k_{*}r)\right|,\qquad\sigma_{Y}=|\lambda|\mathcal{A}_{\mathcal{R}}^{1/2}\left|j_{0}(k_{*}r)\right|,$ (72)
@@
and $\varrho=\mathrm{sgn}(-\lambda j_{0}(k_{*}r)j_{0}^{\prime}(k_{*}r))$. For simplicity, we choose $r_{m}$ to be the same as the Gaussian case, $k_{*}r_{m}\approx 2.74$, which gives $\varrho=\mathrm{sgn}(\lambda)$. With $\varrho^{2}=1$, the PDF (65) reduces to one-dimensional, as $X$ and $Y$ are fully correlated. Together with $X=\mathscr{C}_{\ell}Y$, we have
$X=\frac{\sigma_{X}\mathscr{C}_{\ell}}{\sigma_{X}-\mathrm{sgn}(\lambda)\sigma_{Y}\mathscr{C}_{\ell}},\qquad Y=\frac{\sigma_{X}}{\sigma_{X}-\mathrm{sgn}(\lambda)\sigma_{Y}\mathscr{C}_{\ell}}.$ (73)
And the degenerated PDF of $\mathscr{C}_{\ell}$ becomes
@@
$\mathbb{P}(\mathscr{C}_{\ell})=\frac{1}{\sqrt{2\pi\sigma^{2}_{X}}}\left(\frac{\sigma_{X}}{\sigma_{X}-\mathrm{sgn}(\lambda)\sigma_{Y}\mathscr{C}_{\ell}}\right)^{2}\exp\left[-\frac{1}{2}\left(\frac{\mathscr{C}_{\ell}}{\sigma_{X}-\mathrm{sgn}(\lambda)\sigma_{Y}\mathscr{C}_{\ell}}\right)^{2}\right].$ (74)
@@
Apparently for $\lambda\to 0$, it degenerates to the Gaussian PDF. Roughly speaking, the variance of PDF (74) increases for a negative $\lambda$ (i.e. positive nominal $f_{\mathrm{NL}}$), which enhances PBH formation. As $\mathscr{C}_{\ell}$ increases, the PDF (74) finally approaches the Cauchy distribution $\mathbb{P}(\mathscr{C}_{\ell})\propto\mathscr{C}_{\ell}^{-2}$, which, however, is far from the Type-I-PBH formation region of $30.872<\mathscr{C}_{\ell}<4/$ 3 Biagetti et al. (2021). We calculated the PBH mass function $f_{\mathrm{PBH}}(M)$ for such a monochromatic power spectrum (71) in the ultra-slow-roll inflation ($\lambda=-3$), shown in the left panel of Fig. 6. We choose a typical PBH peak mass of $M_{k^{*}}=10^{-12}M_{\odot}$, and normalize the amplitude of the power spectrum $\mathcal{A_{R}}$ to be $4.151077\times 10^{-3}$, such that $f_{\mathrm{PBH}}=1$, i.e. PBHs can be all the dark matter. Interestingly, as is clearly seen in Fig. 6, for the narrow peak case, the critical infrared scaling $f_{\mathrm{PBH}}(M)\propto M^{1+1/\gamma}\sim M^{3.78}$ which originates from the universal critical behavior is almost parallel to the constraints from the 21-cm signal which is proportional to $M^{4}$ Mittal et al. (2022). So the entire asteroid-mass window is open for narrow-peak power spectrum even with non-Gaussianity.
Another example is the perturbative series (21). Similarly we can define
$\mathscr{C}_{\ell}=XZ,\quad X\equiv-\frac{4}{3}r\mathcal{R}^{\prime}_{\mathrm{g}}(r),\quad Z=1+\frac{6}{5}f_{\mathrm{NL}}\mathcal{R}_{\mathrm{g}}.$ (75)
By the same method, for a narrow spectrum, we can get the PBH abundance $f_{\mathrm{PBH}}$ as a function of $\mathcal{A_{R}}$ for different $f_{\mathrm{NL}}$’s, given by (20) with the PDF of $\mathscr{C}_{\ell}$
$\mathbb{P}\left(\mathscr{C}_{\ell}\right)=\sum_{i=+,-}\frac{|Z_{i}|}{\sqrt{2\pi}|\sigma_{X}Z_{i}^{2}+\mathrm{sgn}(f_{\mathrm{NL}})\mathscr{C}_{\ell}\sigma_{Z}|}\exp\left(-\frac{\mathscr{C}_{\ell}^{2}}{2\sigma_{X}^{2}Z_{i}^{2}}\right),$ (76)
where $\sigma_{Z}$ is given by (72), and
$Z_{\pm}=\frac{1}{2}\pm\sqrt{\frac{1}{4}+\mathrm{sgn}(f_{\mathrm{NL}})\frac{\sigma_{Z}}{\sigma_{X}}\mathscr{C}_{\ell}},\qquad\sigma_{Z}=\frac{6}{5}\mathcal{A}_{\mathcal{R}}^{1/2}\left|f_{\mathrm{NL}}j_{0}(k_{*}r)\right|.$ (77)
When $f_{\mathrm{NL}}\to 0$, the “$+$” branch goes back to Gaussian PDF, while the “--” branch disappears. Roughly speaking, for a positive $f_{\mathrm{NL}}$, the variance of $\mathbb{P}(\mathscr{C}_{\ell})$ increases, which enhances PBH production. In the large-$f_{\mathrm{NL}}$ limit, the PDF reduces to $\chi^{2}$-distribution (a special form of gamma distribution):
$\mathbb{P}(\mathscr{C}_{\ell})\approx\frac{1}{2\sqrt{2\pi\sigma_{X}\sigma_{Z}\mathscr{C}_{\ell}}}\exp\left(-\frac{\mathscr{C}_{\ell}}{2\sigma_{X}\sigma_{Z}}\right).$ (78)
Besides its form, the key point is that in this limit only the product $\sigma_{X}\sigma_{Z}\propto\mathcal{A_{R}}f_{\mathrm{NL}}$ is relevant, and numerical integral shows that $f_{\mathrm{PBH}}=1$ requires $\mathcal{A_{R}}f_{\mathrm{NL}}\sim 0.07$.
@@
We can calculate the total abundance $f_{\mathrm{PBH}}$ as a function of $\mathcal{A_{R}}$ following (20), which is shown in the right panel of Fig. 6, together with the result from ultra-slow-roll inflation and curvaton scenario with $r\to 1$. As a positive $f_{\mathrm{NL}}$ enhances the PBH abundance, the $f_{\mathrm{PBH}}(\mathcal{A_{R}})$ curve moves upward for positive $f_{\mathrm{NL}}$’s, which makes the required $\mathcal{A_{R}}$ for $f_{\mathrm{PBH}}=1$ smaller, as we described in the introduction. Also, because of the higher order terms, the logarithmic relation (57) always generates more PBHs than the quadratic expansion with the same $f_{\mathrm{NL}}$, which requires a slightly smaller $\mathcal{A_{R}}$ for the same abundance of PBH. This is clearly shown for two cases: $f_{\mathrm{NL}}=5/2$ with ultra-slor-roll inflation, and $f_{\mathrm{NL}}=-5/4$ with curvaton scenario with $r\to 1$. The values of $\mathcal{A_{R}}$ required to realize PBH dark matter for $M_{k_{*}}=10^{-12}M_{\odot}$ are listed in Table 1, which can be well fitted by
$\mathcal{A_{R}}\approx\frac{6.6\times 10^{-3}}{1+0.1f_{\mathrm{NL}}},$ (79)
which recovers $\mathcal{A_{R}}f_{\mathrm{NL}}\to 0.07$ when $f_{\mathrm{NL}}\gg 1$.
@@
Figure 6: Left: The PBH mass function $f_{\mathrm{PBH}}(M)$ given by (70) with a monochromatic power spectrum (71) in a ultra-slow-roll inflation with a sharp end, together with the some observational constraints. From left to right: extra-galactic $\gamma$-rays Carr et al. (2020), the 511 keV line from galactic center DeRocco and Graham (2019); Laha (2019), 21-cm signal from EDGES Mittal et al. (2022), Subaru HSC Niikura et al. (2019), EROS Tisserand et al. (2007), OGLE Mroz et al. (2024a, b), and CMB anisotropy with disk accretion Serpico et al. (2020). The total power and central mass are $\mathcal{A_{R}}=4.151\times 10^{-3}$ and $M_{k_{*}}=10^{-12}~{}M_{\odot}$, such that $\int f_{\mathrm{PBH}}(M)\mathrm{d}\ln M=1$. Right: The total PBH abundance $f_{\mathrm{PBH}}$ as a function of the total power $\mathcal{A_{R}}$, for a monochromatic power spectrum (71). From right to left, the curves are for $f_{\mathrm{NL}}=-5/4$ (black dotted), curvaton with $r\to 1$ (gray), Gaussian case ($f_{\mathrm{NL}}=0$) (black thick), $f_{\mathrm{NL}}=1$ (purple dotted), $f_{\mathrm{NL}}=5/2$ (blue dashed), ultra-slow-roll inflation (cyan), $f_{\mathrm{NL}}=10$ (brown dot-dashed), $f_{\mathrm{NL}}=10^{2}$ (magenta dashed), and $f_{\mathrm{NL}}=10^{3}$ (orange), respectively.
@@
0.1 $6.498845\times 10^{-3}$
1 $5.519888\times 10^{-3}$
$f_{\mathrm{NL}}$ $\mathcal{A_{R}}$
5/2 $4.479141\times 10^{-3}$
USR $4.151077\times 10^{-3}$
10 $2.456822\times 10^{-3}$
$10^{2}$ $4.618021\times 10^{-4}$
$10^{3}$ $5.690648\times 10^{-5}$
Table 1: The total power $\mathcal{A_{R}}$ for a monochromatic power spectrum (71) needed to make PBH to be all the dark matter at $M_{k_{*}}=2\times 10^{21}~{}\text{g}$, with different $f_{\mathrm{NL}}$’s. “$r\to 1$” and “USR” stand for curvaton model with $r\to 1$ and ultra-slow-roll inflation, described by a logarithmic $\mathcal{R}(\mathcal{R}_{\mathrm{g}})$ relation (57) with $\lambda=3/2$ ($f_{\mathrm{NL}}=-5/4$) and $\lambda=-3$ ($f_{\mathrm{NL}}=5/2$), respectively.